# Evaluating machine learning models for predicting pesticides toxicity to honey bees

### v2

@@ 5 Conclusions
{acknowledgement}
We thank Wojciech Czech for support and manuscript review. In addition, we thank Piotr Ludynia, Michał Stefanik, and Adam Staniszewski for help with the development of scikit fingerprints and the implementation of functionalities related to this paper. We also thank Alexandra Elbakyan for her work and support for accessibility of science.
The research presented in this paper was financed from funds assigned by the Polish Ministry of Science and Higher Education to the AGH University of Krakow and the Institute of Biochemistry and Biophysics of the Polish Academy of Sciences. Research project was supported by program "Excellence initiative – research university" for the AGH University of Krakow.
Appendix
Hyperparameters and model details
5.0.1 Baselines
@@ 5.0.2 Fingerprints
Fingerprints “Autocorrelation”, Mordred and VSA had no further hyperparameters. For VSA, the full space of possible features was used.
@@
For RDKit2DDescriptors fingerprint, i.e. all RDKit descriptors combined, the only hyperparameter was whether to use CDF normalization proposed by D-MPNN 108.
@@ 5.0.3 Graph kernels
5.0.4 Graph neural networks
@@
5.0.5 Pretrained neural networks
@@
For ChemBERTa, we checked 3 variants available, based on amount of pretraining data: small (“DeepChem/ChemBERTa-5M-MTR” on HuggingFace), medium (“DeepChem/ChemBERTa-10M-MTR”), large (“DeepChem/ChemBERTa-77M-MTR”). For GROVER, we checked base and large versions, as defined by the original paper 74. For MAT, we checked three size variants: small (“mat_masking_200k”), medium (“mat_masking_2M”), large (“mat_masking_20M”).
Full results of molecular fingerprints
@@
Table 5: Classification results on maxmin split.
@@
Full results of pretrained neural models
@@
Fingerprint MCC AUROC Precision Recall
ChemBERTa small 0.25 +- 0.00 71.59% +- 0.00% 34.29% +- 0.00% 57.14% +- 0.00%
ChemBERTa medium 0.37 +- 0.00 74.46% +- 0.00% 42.86% +- 0.00% 64.29% +- 0.00%
ChemBERTa large 0.29 +- 0.00 71.53% +- 0.00% 37.50% +- 0.00% 57.14% +- 0.00%
GROVER base 0.22 +- 0.00 71.46% +- 0.00% 32.86% +- 0.00% 54.76% +- 0.00%
GROVER large 0.21 +- 0.00 66.98% +- 0.00% 31.94% +- 0.00% 54.76% +- 0.00%
MAT small 0.36 +- 0.00 72.29% +- 0.00% 40.58% +- 0.00% 66.67% +- 0.00%
MAT medium 0.18 +- 0.00 66.39% +- 0.00% 30.14% +- 0.00% 52.38% +- 0.00%
MAT large 0.33 +- 0.00 73.17% +- 0.00% 40.32% +- 0.00% 59.52% +- 0.00%
R-MAT 0.31 +- 0.00 70.46% +- 0.00% 36.99% +- 0.00% 64.29% +- 0.00%
Mol2Vec 0.34 +- 0.00 76.57% +- 0.00% 40.62% +- 0.00% 61.90% +- 0.00%

### v3

@@ 5 Conclusions
Acknowledgements
We thank Wojciech Czech for support and useful discussions. In addition, we thank Piotr Ludynia, Michał Stefanik, and Adam Staniszewski for help with the development of scikit-fingerprints and the implementation of functionalities related to this paper. We also thank Alexandra Elbakyan for her work and support for accessibility of science.
The research presented in this paper was financed from funds assigned by the Polish Ministry of Science and Higher Education to the AGH University of Krakow and the Institute of Biochemistry and Biophysics of the Polish Academy of Sciences. Research project was supported by program “Excellence initiative – research university” for the AGH University of Krakow.
Appendix A Hyperparameters and model details
A.1 Baselines
@@ A.2 Fingerprints
Fingerprints Autocorrelation, Mordred and VSA had no further hyperparameters. For VSA, the full space of possible features was used.
@@
For RDKit2DDescriptors fingerprint, i.e. all RDKit descriptors combined, the only hyperparameter was whether to use CDF normalization proposed by D-MPNN [108].
@@ A.3 Graph kernels
A.4 Graph neural networks
@@
A.5 Pretrained neural networks
@@
For ChemBERTa, we checked 3 variants available, based on amount of pretraining data:
  • 1.
    small (“DeepChem/ChemBERTa-5M-MTR” on HuggingFace)
  • 2.
    medium (“DeepChem/ChemBERTa-10M-MTR”)
  • 3.
    large (“DeepChem/ChemBERTa-77M-MTR”)
For GROVER, we checked base and large versions, as defined by the original paper [74].
For MAT, we checked three size variants:
  • 1.
    small (“mat_masking_200k”)
  • 2.
    medium (“mat_masking_2M”)
  • 3.
    large (“mat_masking_20M”)
For R-MAT, we checked variants without and with additional RDKit descriptors.
Appendix B Full results of molecular fingerprints
@@
Table 4: Classification results on time split.
@@
Table 5: Classification results on maxmin split.
Appendix C Full results of pretrained neural models
@@
Table 6: Classification results on time split.
Fingerprint MCC AUROC Precision Recall
ChemBERTa small 0.25 $\pm$ 0.00 71.59% $\pm$ 0.00% 34.29% $\pm$ 0.00% 57.14% $\pm$ 0.00%
ChemBERTa medium 0.37 $\pm$ 0.00 74.46% $\pm$ 0.00% 42.86% $\pm$ 0.00% 64.29% $\pm$ 0.00%
ChemBERTa large 0.29 $\pm$ 0.00 71.53% $\pm$ 0.00% 37.50% $\pm$ 0.00% 57.14% $\pm$ 0.00%
GROVER base 0.22 $\pm$ 0.00 71.46% $\pm$ 0.00% 32.86% $\pm$ 0.00% 54.76% $\pm$ 0.00%
GROVER large 0.21 $\pm$ 0.00 66.98% $\pm$ 0.00% 31.94% $\pm$ 0.00% 54.76% $\pm$ 0.00%
MAT small 0.36 $\pm$ 0.00 72.29% $\pm$ 0.00% 40.58% $\pm$ 0.00% 66.67% $\pm$ 0.00%
MAT medium 0.18 $\pm$ 0.00 66.39% $\pm$ 0.00% 30.14% $\pm$ 0.00% 52.38% $\pm$ 0.00%
MAT large 0.33 $\pm$ 0.00 73.17% $\pm$ 0.00% 40.32% $\pm$ 0.00% 59.52% $\pm$ 0.00%
R-MAT 0.31 $\pm$ 0.00 70.46% $\pm$ 0.00% 36.99% $\pm$ 0.00% 64.29% $\pm$ 0.00%
Mol2Vec 0.34 $\pm$ 0.00 76.57% $\pm$ 0.00% 40.62% $\pm$ 0.00% 61.90% $\pm$ 0.00%